If shuanais now 4 years older than her sister. Two years ago, shuanais was three times as old as the sister was. How old is each girl?

Answer by fractalier(6550)

Call her and her sister's age NOW, x and y. Then we write
x = y + 4
Two years ago
x - 2 = 3(y - 2)
Now substitute the first equation into the second equation...we get
y+4 - 2 = 3(y - 2)
y + 2 = 3y - 6
8 = 2y and
y = 4
so that
x = 8